OpenStudy (anonymous):

solve for x: (2x-4)^2=1 A. 2,3 B. 3/2, 5/2 C. 3/2, 3/5 D.3/2, -5/2

OpenStudy (anonymous):

x=(sqrt(1)+4)/2

OpenStudy (toxicsugar22):

hi can you help me. I t is posted in chemiisty

OpenStudy (johnweldon1993):

from the looks of it, they want you to use the quadratic formula (2x-4)^2=1 multiply that out 4x² - 8x + 16 = 1 subtract 1 from both sides 4x² - 8x + 15 = 0 now use the quadratic formula...this will give you 2 results...which IS what you are looking for \[\frac{ -b \pm \sqrt{b² - 4ac} }{ 2a }\] in your case ... a = 4 b=-8 c = 15 just plug those number into that equation...and you'll get your answers